Fermi National Laboratory
 
Operation Summary for November 28, 2011 - November 30, 2011
 

Monday November 28

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), with NuMI off for a PLC replacement, and with the MiniBooNE taking beam.

During the day shift, Operations established beam to NuMI at 9:43 AM and then rebooted a NuMI BPM (NBPM6N). Operations established beam to the Meson FTBFT-978 experiment at 11:10 AM.

During the evening shift, at 4:27 PM, the NuMI beam permit tripped off due to an MI incomplete RF gap for the first of two times in about an hour, but no RF stations had tripped off; it reset both times. Operations disabled the T-978 beam around 7:10 PM. At 11:52 PM, the NuMI beam permit tripped off due to a temperature problem with a string of dipole magnets (E:HV101); it would not immediately reset, but did five minutes later. Operators contacted a NuMI person.

Tuesday November 29

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE.

During the midnight shift, at 4:11 AM, the NuMI beam permit tripped off due to E:HV101 magnet temperature; operators contacted a NuMI person, lowered the rep rate, and reset the power supply.

During the day shift, Operations established beam to T-978 at 8:23 AM. Operations set up HV101 to see if it would trip on a magnet over temperature at 9:38 AM; it tripped around 10:22 AM and operators prepared MI for an access to inspect the magnet string. Operator prepared Booster for an access at 10:31 AM. Operations prepared the Muon PreVault area for a one-hour access at 10:53 AM. RF personnel reported at 11:51 AM that the power amplifier for a MI RF station (MIRF15) needed to be replaced; they accessed, but found a LCW leak and began repairs. At 12:46 PM, personnel investigating the HV101 magnet string reported that they would back flush the magnets. Operators prepared NuMI for access at 1:51 PM.

During the evening shift, Operations established beam to the Booster dump at 4:14 PM. Operators turned on MI at 5:08 PM. Operations established beam to MiniBooNE at 5:26 PM. Operations reestablished beam to T-978 at 5:56 PM, but only for an hour. Operations established beam to NuMI around 6:45 PM. Operators reported problems with a Booster RF station (BRF16) at 8:01 PM; operators called in a RF person to investigate.

Wednesday November 30

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE.

During the midnight shift, at 2:01 AM, a NuMI BPM (NBPM6N) stopped responding; operators rebooted the BPM. At 2:27 AM, NBPM65 stopped responding; operators rebooted the BPM. At 3:21 AM, BRF13 tripped off and would not reset; operators bypassed the station, lowered the NuMI beam intensity, and then tuned the Booster to gain back some intensity. Operators contacted RF personnel.

Operations Coordinator (OpsCo) Plans for Today and the Holiday Weekend

The OpsCo approved of the plans by Booster, MI, and Muon personnel to conduct studies. Booster reported that BRF13 needs a tunnel access to repair; they will wait for an opportunity to arise.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/30/2011    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory